# The evolution of the galaxy stellar mass function and star formation rates in the COLIBRE simulations from redshift 17 to 0

### v1

@@ 3.1 Evolution of the galaxy stellar mass function
Fig. 1 shows the evolution of the GSMF in the colibre simulations from $z=0$ to $z=4$, and Fig. 2 extends the evolution to higher redshifts, from $z=5$ to z=17z=17444The choice of $z=17$ is motivated by the fact that this is the highest integer redshift for which our lowest-resolution simulation (m7) contains a sufficient number of subhaloes with non-zero stellar masses ($\approx 60$), so that the GSMF can still be estimated.. The GSMF is computed using fixed mass bins of $0.2$ dex. The colibre models with thermal AGN feedback at m7, m6, and m5 resolutions are depicted by the solid dark-red, orange, and light-blue curves, respectively. To highlight limitations due to resolution and volume, the curves transition from solid to dotted at stellar masses below $10$ times the mean initial baryonic particle mass (i.e. where galaxies contain fewer than $\sim 10$ stellar particles) or where fewer than 10 galaxies are present per bin. We emphasise that the chosen resolution limit is shown only to guide the eye and does not necessarily correspond to the true resolution limit. The binned GSMF values for the L200m6 simulation at $0<z<17$ are provided in Appendix E, and those for other simulations will be made available on the colibre website upon publication [^5] https://colibre-simulations.org/.
@@
  • The GSMF from Harvey et al. (2025) based on JWST/NIRCam imaging of $1,120$ galaxies at $6.5<z<13.5$ from multiple JWST fields, including CEERS, JADES, and PEARLS (covering a total area of $\approx 187$ arcmin2). Stellar masses were estimated using bagpipes, adopting the Bruzual & Charlot (2003) stellar population models with log-normal SFHs. The allowed metallicity values ranged from $0.005\,\rm Z_{\odot}$ to $5\,\rm Z_{\odot}$. Photometric redshifts were determined using eazy.
    Figure 3: Evolution of the stellar-to-halo mass relation (SHMR) for central subhaloes in the colibre simulations at m7 (left), m6 (middle), and m5 (right) resolutions. Solid curves represent the median stellar-to-halo mass ratios from the simulations, while the shaded regions show the 16th to 84th percentile scatter. Different colours indicate different redshifts, ranging from $z=0.1$ (dark-blue) to $z=11$ (yellow). The solid lines switch to dotted at halo masses corresponding to galaxies with $M_{*}<10$ times the mean initial baryonic particle mass, or where fewer than 10 objects fall within a given halo mass bin. For comparison, the semi-empirical SHMRs from Moster et al. (2018) and Behroozi et al. (2019) at $z=0.1$ are shown by black symbols. Here we do not add a lognormal scatter to galaxy stellar masses from the simulations because the Eddington bias is absent in the comparison data. The colibre SHMR is consistent with the semi-empirical comparison data at $M_{\rm halo}\gtrsim 10^{11}~\mathrm{M_{\odot}}$, where it shows excellent convergence with resolution, peaks at a halo mass $M_{\rm halo}\sim 10^{12}~\mathrm{M_{\odot}}$, and exhibits only weak evolution with redshift.
@@
It is evident from Figs. 1 and 2 that the colibre simulations are in good agreement with the observationally inferred GSMF at all redshifts where observational data are available. At $z=0$, the excellent agreement of the colibre m7 model is expected, as it was calibrated using emulators to match the GSMF from Driver et al. (2022). Although the m6 and m5 models were calibrated manually, they also show excellent agreement, with deviations within $\approx 0.1$ dex and no systematic offset from the observations.
For all colibre resolutions, deviations from the data increase with redshift, peaking at $z\approx 2-4$, where the simulations predict a GSMF with number densities that are on average $\approx 0.2-0.3$ dex higher than observed at $M_{*}\lesssim 10^{10.5}~\mathrm{M_{\odot}}$. The smallest discrepancies (around $0.2$ dex) are found in the m7 and m6 models when compared to the measurements from Leja et al. (2020), whereas the largest deviations – up to $\approx 0.4$ dex – occur in the m5 model relative to Weaver et al. (2023b) and Shuntov et al. (2025). We emphasize, however, that multiple studies have shown that, depending on the assumed SFH and SPS models, inferred stellar masses at $z\approx 2-4$ may be systematically offset by 0.10.30.1-0.3 dex relative to the true values (e.g. Katsianis et al., 2020; Shuntov et al., 2025). With this in mind, reducing the stellar masses in the simulations by $0.2$ dex at $M_{*}\lesssim 10^{10.5}~\mathrm{M_{\odot}}$ would bring the simulations into much closer agreement with the data (within $\approx 0.15$ dex).
@@
At the highest redshifts shown in Fig. 2, from $z=12$ to $17$, colibre predicts a gradual evolution of the GSMF without sharp drop-offs, with the normalization changing by $\approx 2$ dex over this redshift interval. At $z=12$, we compare to the only available observational dataset, Harvey et al. (2025), who provide GSMF estimates based on JWST/NIRCam imaging at a median redshift of $z\approx 11.94$ and in two stellar mass bins: $M_{*}\sim 10^{8.125}$ and $10^{8.875}~\mathrm{M_{\odot}}$, each containing 5 objects (see their table 3). We find that colibre is consistent with the lower mass bin within $1\sigma$ but underpredicts the value in the higher mass bin by $\approx 2\sigma$. Note, however, that we are not accounting for potential systematic uncertainties. A shift in $M_{*}$ by $0.3$ dex, which is a relatively conservative estimate for the magnitude of the systematic error on stellar mass at $z=12$, would bring the colibre prediction into agreement with the second bin within $1\sigma$. Alternatively, increasing the assumed stellar mass uncertainty for simulated galaxies from our $z\geq 2$ default (and likely conservative for $z=12$) value of $0.3$ dex to $0.5$ dex would also resolve the discrepancy (see Fig. 20 in Appendix D).
@@
Figure 4: Top panel: Evolution of the cosmic stellar mass density (CSMD) in the colibre m7, m6, and m5 simulations. Different resolutions are indicated by different colours. We show the total CSMD, including all stellar mass within the simulated volumes (solid lines), and the CSMD contributed only by galaxies with stellar masses $M_{*}\geq 10^{8}\,\mathrm{M_{\odot}}$ (dash-dotted lines), where the stellar mass is measured within 3D spherical apertures of radius 50 pkpc. While the total CSMD reflects the ‘raw’ simulation values, the CSMD shown by the dash-dotted lines incorporates common observational cuts and can be directly compared to observations. For clarity, the solid and dash-dotted lines for the m7 and m5 simulations are truncated at $z<4$ due to their overlap with the m6 lines. For comparison, a compilation of recent observational measurements is shown as black symbols. No lognormal scatter is added to the simulated galaxy stellar masses, as the comparison data account for Eddington bias when estimating the CSMD. Bottom panel: Relative difference in the CSMD between simulations at different resolutions, shown for both versions of CSMD and using the m6 simulation as the baseline. The CSMD in colibre agrees well with the observational data at all redshifts and shows very good convergence between resolutions for $z<8$. Selecting only galaxies with $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ reduces the CSMD by more than $0.5$ dex at $z\gtrsim 6$, while having only a minor impact below $z=2$.
@@ 3.3 Cosmic stellar mass density
For clarity, the solid and dash-dotted lines for the m5 and m7 simulations are truncated below $z=4$, where they overlap with the CSMD from the m6 simulation. The bottom panel shows the relative difference in CSMD between the m7 and m5 simulations compared to the m6 simulation. The relative difference for a simulation at resolution mXX, where $X$ represents 5, 6, or 7, is calculated as $\Delta\mathrm{CSMD}_{\mathrm{m}X}\equiv\mathrm{CSMD}_{\mathrm{m}X}/\mathrm{CSMD}_{\mathrm{m6}}-1$, where $\mathrm{CSMD}_{\mathrm{m}X}$} is the CSMD from the mXX simulation. We compute these differences for two cases: using all stellar mass in the simulated volume (solid lines) and applying the $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ selection criterion (dash-dotted lines). Because the L50m5 simulation has not yet reached $z=0$ at the time of writing (it is currently at $z=2.5$; see Table 1), we show the CSMD at m5 resolution from the $50^{3}$ cMpc3 volume for $z>2.5$, and from the $25^{3}$ cMpc3 volume at lower redshifts. In Appendix A, we demonstrate that at $z<3$, the CSMDs in these two volumes converge very well, which justifies the transition from the $50^{3}$ cMpc3 to $25^{3}$ cMpc3 volume at $z=2.5$.
For comparison, we include CSMD estimates from the same nine studies shown in Figs. 1 and 2, as well as additional measurements from Driver et al. (2018). Specifically, we show the $z=0$ CSMD value from Driver et al. (2022), obtained by integrating the GAMA DR4 GSMF down to $M_{*}=10^{6.75}~\mathrm{M_{\odot}}$. This estimate accounts for re-normalization of the GAMA fields to a larger region selected from the SDSS DR16, along with a correction for CSMD evolution between the GAMA sample’s median observed redshift ($z\approx 0.079$) and $z=0$. We also include CSMD estimates from Leja et al. (2020), Stefanon et al. (2021), Thorne et al. (2021), Weaver et al. (2023b), Weibel et al. (2024), Shuntov et al. (2025), Navarro-Carrera et al. (2024), and Harvey et al. (2025), all derived by integrating their respective best-fitting Schechter functions down to $M_{*}=10^{8}~\mathrm{M_{\odot}}$. Finally, we include the $0<z<5$ CSMD estimates from Driver et al. (2018), based on a combined dataset of $\sim 10^{6}$ galaxies from GAMA DR2, G10-COSMOS, and the 3D-HST survey. In their analysis, the authors fit a spline to their GSMF measurements and integrated it down to $M_{*}\sim 10^{7}~\mathrm{M_{\odot}}$, relying increasingly on extrapolation at higher redshifts.
@@
Next, by comparing the observational data with the dash-dotted lines, which best mimic the observational selection criteria, we find that the CSMD in colibre is generally in good agreement with the observed CSMD across all redshifts. Specifically, the colibre CSMD at $z=0$ best matches the measurements from Driver et al. (2022); from $z=0$ to $z=1$, it is in good agreement with Thorne et al. (2021); between $z=1$ and $4$, colibre remains consistent with the CSMD from Driver et al. (2018) and Leja et al. (2020); and at $4<z<12$, it agrees reasonably well with the CSMD reported by Shuntov et al. (2025). At the same time, colibre slightly overestimates the CSMD reported by Weaver et al. (2023b) and Shuntov et al. (2025) at $1.5<z<4$, and underestimates the measurements from Navarro-Carrera et al. (2024) at $z>6$ and from Weibel et al. (2024) at $z>7$. These discrepancies are consistent with those seen in the GSMF (see Figs. 1 and 2, and §3.1 for discussion). The colibre CSMD is consistent with the observations from Stefanon et al. (2021) and Harvey et al. (2025) over the redshift range $6<z<9$, but at $z\approx 10$ it falls $\approx 0.5$ dex below the measurement from Harvey et al. (2025), while exceeding that of Stefanon et al. (2021) by $\approx 1$ dex.
Finally, in the lower panel, we observe that the relative differences in the CSMD between the m5 and m6 simulations and between the m6 and m7 simulations remain within $\approx 20$ per cent at $0<z<8$, but increase sharply beyond $40$ per cent at higher redshifts. As we will clarify in Appendix A, this sharp increase in the relative differences at $z>8$ is caused by resolution effects.
@@ 3.4 Cosmic star formation rate density
The top panel of Fig. 5 shows the evolution of the cosmic star formation rate density (CSFRD). The solid curves show the SFR from all star-forming gas particles in the simulation, while the dash-dotted curves indicate the SFR from gas particles that are bound to and located within 50 pkpc of subhaloes containing at least $10^{7}~\mathrm{M_{\odot}}$ of stellar mass. This mass threshold is chosen to facilitate a more consistent comparison with JWST observations at high redshift. As in Fig. 4, we highlight the m6 simulation, while the CSFRD curves for the m7 and m5 simulations are truncated below $z=4$ for visual clarity. The bottom panel displays the differences in CSFRD relative to the m6 simulation. As in Fig. 4, we show the predictions of the colibre m5 model from the $50^{3}$ cMpc3 volume for $z>2.5$, and from the $25^{3}$ cMpc3 volume at lower redshifts, since the larger volume simulation is not yet available at $z<2.5$.
Figure 5: Evolution of the cosmic star formation rate density (CSFRD). The figure follows the same layout as Fig. 4. We show the total CSFRD, contributed by all star-forming gas particles within the simulated volumes (solid lines), and the CSFRD from galaxies with stellar masses $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ (dash-dotted lines). Black symbols represent a compilation of observational data from Robotham & Driver (2011); Driver et al. (2012); Novak et al. (2017); Driver et al. (2018); Gruppioni et al. (2020); Khusanova et al. (2021); Enia et al. (2022); Cochrane et al. (2023); Bouwens et al. (2023); Harikane et al. (2023); Donnan et al. (2024); Covelo-Paz et al. (2025); Fu et al. (2025); Weibel et al. (2025a). The CSFRD in colibre exhibits a plausible shape and normalization, showing good agreement with the data at all redshifts and very good numerical convergence at $z<8$.
@@
The colibre simulations exhibit good numerical convergence for both particle selection criteria. At $0<z<8$, the relative differences between the simulations remain within $\approx 20$ per cent. At higher redshifts ($z\gtrsim 10$), the convergence worsens, with the higher-resolution simulations predicting systematically higher SFRs than their lower-resolution counterparts. As clarified later in Fig. 6, this behaviour is driven by resolution effects. At $z\gtrsim 10$, the CSFRD is dominated by SFRs from low-mass subhaloes ($M_{\rm halo}\lesssim 10^{9.5}~\mathrm{M_{\odot}}$). Higher-resolution simulations resolve more of these low-mass objects and begin to track them at earlier times, which together leads to a higher CSFRD.
@@
Turning to the comparison with observational data, at $z\approx 0$, colibre predicts a CSFRD of $\approx 0.012~\mathrm{M_{\odot}\,yr^{-1}~cMpc^{-3}}$, which is consistent with the $z\sim 0$ measurements from Driver et al. (2012), Robotham & Driver (2011), and Driver et al. (2018). From $z=0$ to $2$, the CSFRD in the simulations increases by about $0.7$ dex, closely following the measurements from Novak et al. (2017) and Driver et al. (2018). The simulated CSFRD peaks at $\approx 0.075~\mathrm{M_{\odot}\,yr^{-1}\,cMpc^{-3}}$ between $z=2$ and $3$, in reasonable agreement with Novak et al. (2017) and Driver et al. (2018). However, it falls short by 0.20.40.2-0.4 dex compared to the measurements of Gruppioni et al. (2020), Enia et al. (2022), and Cochrane et al. (2023). The factor of $\approx 2$ discrepancy at $z\approx 2$ between the CSFRD computed from intrinsic SFRs in simulations (or inferred from the observed CSMD) and that derived from direct SFR indicators in observations has long been recognized (e.g. Wilkins et al., 2008; Yu & Wang, 2016; Wilkins et al., 2019; Shuntov et al., 2025). The source of this discrepancy remains unclear, but it is often attributed to systematic offsets in observational data, arising from uncertainties in the calibration of SFR indicators and assumptions regarding the dust attenuation curve, the stellar IMF, and/or gas metallicity.

### v2

@@ 3.1 Evolution of the galaxy stellar mass function
Fig. 1 shows the evolution of the GSMF in the colibre simulations from $z=0$ to $z=4$, and Fig. 2 extends the evolution to higher redshifts, from $z=5$ to z=17z=17555The choice of $z=17$ is motivated by the fact that this is the highest integer redshift for which our lowest-resolution simulation (m7) contains a sufficient number of subhaloes with non-zero stellar masses ($\approx 60$), so that the GSMF can still be estimated.. The GSMF is computed using fixed mass bins of $0.2$ dex. The colibre models with thermal AGN feedback at m7, m6, and m5 resolutions are depicted by the solid dark-red, orange, and light-blue curves, respectively. To highlight limitations due to resolution and volume, the curves transition from solid to dotted at stellar masses below $10$ times the mean initial baryonic particle mass (i.e. where galaxies contain fewer than $\sim 10$ stellar particles) or where fewer than 10 galaxies are present per bin. We emphasize that the chosen resolution limit is shown only to guide the eye and does not necessarily correspond to the true resolution limit. The binned GSMF values for the L200m6 simulation at $0<z<17$ are provided in Appendix E, and those for other simulations will be made available on the colibre website upon publication [^6] https://colibre-simulations.org/.
@@
  • The GSMF from Harvey et al. (2025) based on JWST/NIRCam imaging of $1,120$ galaxies at $6.5<z<13.5$ from multiple JWST fields, including CEERS, JADES, and PEARLS (covering a total area of $\approx 187$ arcmin2). Stellar masses were estimated using bagpipes, adopting the Bruzual & Charlot (2003) stellar population models with lognormal SFHs. The allowed metallicity values ranged from $0.005\,\rm Z_{\odot}$ to $5\,\rm Z_{\odot}$. Photometric redshifts were determined using eazy.
  • The GSMF from Euclid Collaboration et al. (2025b) at $0.2<z<6.5$ derived using the Cosmic Dawn Survey pre-launch catalogues (Euclid Collaboration et al., 2025c), with an effective area of $\approx 10.13~\textrm{deg}^{2}$ and $2,091,740$ galaxies in the final sample after all selection criteria were applied. Multi-wavelength data were provided by Spitzer/IRAC, the Hyper Suprime-Cam instrument on the Subaru Telescope, and the MegaCam instrument on the Canada–France–Hawaii Telescope. Photometric redshifts were inferred using LePhare. Stellar masses were also derived with LePhare, using the same configuration and stellar templates as in COSMOS2020 (Weaver et al., 2023b).
    Figure 3: Evolution of the stellar-to-halo mass relation (SHMR) for central subhaloes in the colibre simulations at m7 (left), m6 (middle), and m5 (right) resolutions. Solid curves represent the median stellar-to-halo mass ratios from the simulations, while the shaded regions show the 16th to 84th percentile scatter. Different colours indicate different redshifts, ranging from $z=0.1$ (dark blue) to $z=11$ (yellow). The solid lines switch to dotted at halo masses corresponding to galaxies with $M_{*}<10$ times the mean initial baryonic particle mass, or where fewer than 10 objects fall within a given halo mass bin. For comparison, the semi-empirical SHMRs from Moster et al. (2018) and Behroozi et al. (2019) at $z=0.1$ are shown by black symbols. Here we do not add a lognormal scatter to galaxy stellar masses from the simulations because the Eddington bias is absent in the comparison data. The colibre SHMR is consistent with the semi-empirical comparison data at $M_{\rm halo}\gtrsim 10^{11}~\mathrm{M_{\odot}}$, where it shows excellent convergence with resolution, peaks at a halo mass $M_{\rm halo}\sim 10^{12}~\mathrm{M_{\odot}}$, and exhibits only weak evolution with redshift.
@@
It is evident from Figs. 1 and 2 that the colibre simulations are in good agreement with the observationally inferred GSMF at all redshifts where observational data are available. At $z=0$, the excellent agreement of the colibre m7 model is expected, as it was calibrated using emulators to match the GSMF from Driver et al. (2022). Although the m6 and m5 models were calibrated manually, they also show very good agreement, with deviations within $\approx 0.1$ dex and no systematic offset from the observations.
For all colibre resolutions, deviations from the data increase with redshift, peaking at $z\approx 2-4$, where the simulations predict a GSMF with number densities that are on average $\approx 0.3$ dex higher than observed at $M_{*}\lesssim 10^{10.5}~\mathrm{M_{\odot}}$. The smallest discrepancies (around $0.2$ dex) are found for the m7 and m6 models when compared to the measurements of Leja et al. (2020), whereas the largest deviations – up to $\approx 0.4$ dex – occur in the m5 model relative to Weaver et al. (2023b) and Shuntov et al. (2025), as well as in the m7 and m6 models relative to Euclid Collaboration et al. (2025b). While these minor discrepancies may point to limitations of the simulations, we also note that multiple studies have shown that, depending on the assumed SFH and SPS models, inferred stellar masses at $z\approx 2-4$ may be systematically offset by 0.10.30.1-0.3 dex relative to the true values (e.g. Katsianis et al., 2020; Shuntov et al., 2025). With this in mind, reducing the stellar masses in the simulations by $0.2$ dex at $M_{*}\lesssim 10^{10.5}~\mathrm{M_{\odot}}$ would bring the simulations into much closer agreement with the data, within $\approx 0.2$ dex.
@@
At the highest redshifts shown in Fig. 2, from $z=12$ to $17$, colibre predicts a gradual evolution of the GSMF without sharp drop-offs, with the normalization changing by $\approx 2$ dex over this redshift interval. At $z=12$, we compare to the only available observational dataset, Harvey et al. (2025), who provide GSMF estimates based on JWST/NIRCam imaging at a median redshift of $z\approx 11.94$ and in two stellar mass bins: $M_{*}\sim 10^{8.125}$ and $10^{8.875}~\mathrm{M_{\odot}}$, each containing 5 objects (see their table 3). We find that colibre is consistent with the lower mass bin within $1\sigma$ at all three resolutions but underpredicts the value in the higher mass bin by up to $\approx 2\sigma$, depending on the resolution. Note, however, that we are not accounting for potential systematic uncertainties. A shift in $M_{*}$ by $0.3$ dex, which is a relatively conservative estimate for the magnitude of the systematic error on stellar mass at $z=12$, would bring the colibre predictions at all three resolutions into agreement with the second bin within $1\sigma$. Alternatively, increasing the assumed stellar mass uncertainty for simulated galaxies from our $z\geq 2$ default (and likely conservative for $z=12$) value of $0.3$ dex to $0.5$ dex would also resolve the discrepancy (see Fig. 20 in Appendix D).
@@
Figure 4: Top panel: Evolution of the cosmic stellar mass density (CSMD) in the colibre m7, m6, and m5 simulations. Different resolutions are indicated by different colours. We show the total CSMD, including all stellar mass within the simulated volumes (solid lines), and the CSMD contributed only by galaxies with stellar masses $M_{*}\geq 10^{8}\,\mathrm{M_{\odot}}$ (dash-dotted lines), where the stellar mass is measured within 3D spherical apertures of radius 50 pkpc. While the total CSMD reflects the ‘raw’ simulation values, the CSMD shown by the dash-dotted lines incorporates common observational cuts and can be directly compared to observations. For clarity, the solid and dash-dotted lines for the m7 and m5 simulations are truncated at $z<4$ due to their overlap with the m6 lines. For comparison, a compilation of recent observational measurements is shown as black symbols. No lognormal scatter is added to the simulated galaxy stellar masses, as the comparison data account for Eddington bias when estimating the CSMD. Bottom panel: Relative difference in the CSMD between simulations at different resolutions, shown for both versions of CSMD and using the m6 simulation as the baseline. The CSMD in colibre agrees well with the observational data at all redshifts and shows very good convergence between resolutions for $z<6$. Selecting only galaxies with $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ reduces the CSMD by more than $0.5$ dex at $z\gtrsim 6$, while having only a minor impact below $z=2$.
@@ 3.3 Cosmic stellar mass density
For clarity, the solid and dash-dotted lines for the m5 and m7 simulations are truncated below $z=4$, where they overlap with the CSMD from the m6 simulation. The bottom panel shows the relative difference in CSMD between the m7 and m5 simulations compared to the m6 simulation. The relative difference for a simulation at resolution mXX, where $X$ represents 5, 6, or 7, is calculated as $\Delta\mathrm{CSMD}_{\mathrm{m}X}\equiv\mathrm{CSMD}_{\mathrm{m}X}/\mathrm{CSMD}_{\mathrm{m6}}-1$, where $\mathrm{CSMD}_{\mathrm{m}X}$} is the CSMD from the mXX simulation. We compute these differences for two cases: using all stellar mass in the simulated volume (solid lines) and applying the $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ selection criterion (dash-dotted lines). Because the L100m5 and L50m5 simulations have not yet progressed to $z=0$ at the time of writing (they are currently at $z=4.6$ and $z=0.9$, respectively; see Table 1), we show the CSMD at m5 resolution from the $100^{3}$ cMpc3 volume for $z>4.6$, from $50^{3}$ cMpc3 between $z=0.9$ and $z=4.6$, and from the $25^{3}$ cMpc3 volume at lower redshifts. In Appendix A, we demonstrate that at $z<4$, the CSMDs in these three volumes converge very well, while at higher redshifts the $100^{3}$ cMpc3 volume yields a slightly higher CSMD.
For comparison, we include CSMD estimates from the same ten studies shown in Figs. 1 and 2, as well as additional measurements from Driver et al. (2018). Specifically, we show the $z=0$ CSMD value from Driver et al. (2022), obtained by integrating the GAMA DR4 GSMF down to $M_{*}=10^{6.75}~\mathrm{M_{\odot}}$. This estimate accounts for re-normalization of the GAMA fields to a larger region selected from the SDSS DR16, along with a correction for CSMD evolution between the GAMA sample’s median observed redshift ($z\approx 0.079$) and $z=0$. We also include CSMD estimates from Leja et al. (2020), Stefanon et al. (2021), Thorne et al. (2021), Weaver et al. (2023b), Weibel et al. (2024), Shuntov et al. (2025), Navarro-Carrera et al. (2024), Harvey et al. (2025), and Euclid Collaboration et al. (2025b), all derived by integrating their respective best-fitting Schechter functions down to $M_{*}=10^{8}~\mathrm{M_{\odot}}$. Finally, we include the $0<z<5$ CSMD estimates from Driver et al. (2018), based on a combined dataset of $\sim 10^{6}$ galaxies from GAMA DR2, G10-COSMOS, and the 3D-HST survey. In their analysis, the authors fit a spline to their GSMF measurements and integrated it down to $M_{*}\sim 10^{7}~\mathrm{M_{\odot}}$, relying increasingly on extrapolation at higher redshifts.
@@
Next, by comparing the observational data with the dash-dotted lines, which best mimic the observational selection criteria, we find that the CSMD in colibre is generally in good agreement with the observed CSMD across all redshifts. Specifically, the colibre CSMD at $z=0$ best matches the measurements from Driver et al. (2022); from $z=0$ to $z=1$, it is in good agreement with Thorne et al. (2021); between $z=1$ and $4$, colibre remains consistent with the CSMD from Driver et al. (2018) and Leja et al. (2020); and at $4<z<12$, it agrees reasonably well with the CSMD reported by Shuntov et al. (2025). At the same time, colibre slightly overestimates the CSMD reported by Weaver et al. (2023b), Shuntov et al. (2025), and Euclid Collaboration et al. (2025b) at $1.5<z<4$, and underestimates the measurements from Navarro-Carrera et al. (2024) at $z>6$ and from Weibel et al. (2024) at $z>7$. These discrepancies are consistent with those seen in the GSMF (see Figs. 1 and 2, and §3.1 for discussion). The colibre CSMD is consistent with the observations from Stefanon et al. (2021) and Harvey et al. (2025) over the redshift range $6<z<9$, but at $z\approx 10$ it falls $\approx 0.5$ dex below the measurement from Harvey et al. (2025), while exceeding that of Stefanon et al. (2021) by $\approx 1$ dex.
Finally, in the lower panel, we observe that the relative differences in the CSMD between the m5 and m6 models, and between the m6 and m7 models, remain within $\approx 20$ per cent at $0<z<5$, increase to within $\approx 40$ per cent at $5<z<6$, and rise to beyond $40$ per cent at higher redshifts. As we clarify in Appendix A, the sharp rise in the relative differences at $z>7$ is caused by resolution effects, while the increase in the difference between the m6 and m5 resolutions at $5<z<7$ is due to the switch from the L050m5 simulation to L100m5 at $z=5$, with the larger volume containing a massive protocluster at $z\sim 5-6$ that is atypical for a $(100~\mathrm{cMpc})^{3}$ volume and hence yields a higher CSMD compared to L200m6.
@@ 3.4 Cosmic star formation rate density
The top panel of Fig. 5 shows the evolution of the cosmic star formation rate density (CSFRD). The solid curves show the SFR from all star-forming gas particles in the simulation, while the dash-dotted curves indicate the SFR from gas particles that are bound to and located within 50 pkpc of subhaloes containing at least $10^{7}~\mathrm{M_{\odot}}$ of stellar mass. This mass threshold is chosen to facilitate a more consistent comparison with JWST observations at high redshift. As in Fig. 4, we highlight the m6 simulation, while the CSFRD curves for the m7 and m5 simulations are truncated below $z=4$ for visual clarity. The bottom panel displays the differences in CSFRD relative to the m6 simulation. As in Fig. 4, we show the predictions of the colibre m5 model from the $100^{3}$ cMpc3 volume for $z>4.6$, from the $50^{3}$ cMpc3 volume for $0.9<z<4.6$, and from the $25^{3}$ cMpc3 volume at lower redshifts.
Figure 5: Evolution of the cosmic star formation rate density (CSFRD). The figure follows the same layout as Fig. 4. We show the total CSFRD, contributed by all star-forming gas particles within the simulated volumes (solid lines), and the CSFRD from galaxies with stellar masses $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ (dash-dotted lines). Black symbols represent a compilation of observational data from Robotham & Driver (2011); Driver et al. (2012); Novak et al. (2017); Driver et al. (2018); Gruppioni et al. (2020); Khusanova et al. (2021); Enia et al. (2022); Cochrane et al. (2023); Bouwens et al. (2023); Harikane et al. (2023); Donnan et al. (2024); Covelo-Paz et al. (2025); Fu et al. (2025); Weibel et al. (2025a). The CSFRD in colibre exhibits a plausible shape and normalization, showing good agreement with the data at all redshifts and very good numerical convergence at $z<6$.
@@
The colibre simulations exhibit good numerical convergence for both particle selection criteria. At $0<z<6$, the relative differences between the simulations remain within $\approx 20$ per cent. At $z>6$, the convergence of the CSFRD from $M\geq 10^{7}~\mathrm{M_{\odot}}$ galaxies becomes worse, driven by differences in the simulated volumes used at each resolution (see Appendix A). At even higher redshifts ($z\gtrsim 10$), the predicted total CSFRD begins to strongly diverge, with higher-resolution simulations yielding systematically higher SFRs than their lower-resolution counterparts. As clarified later in Fig. 6, this behaviour is driven by resolution effects. At $z\gtrsim 10$, the CSFRD is dominated by SFRs from low-mass subhaloes ($M_{\rm halo}\lesssim 10^{9.5}~\mathrm{M_{\odot}}$). Higher-resolution simulations resolve more of these low-mass objects and begin to track them at earlier times, which together lead to a higher CSFRD.
@@
Turning to the comparison with observational data, at $z\approx 0$, colibre predicts a CSFRD of $\approx 0.012-0.013~\mathrm{M_{\odot}\,yr^{-1}~cMpc^{-3}}$, which is consistent with the $z\sim 0$ measurements from Driver et al. (2012), Robotham & Driver (2011), and Driver et al. (2018). From $z=0$ to $2$, the CSFRD in the simulations increases by about $0.7$ dex, closely following the measurements from Novak et al. (2017) and Driver et al. (2018). The simulated CSFRD peaks at $\approx 0.075~\mathrm{M_{\odot}\,yr^{-1}\,cMpc^{-3}}$ between $z=2$ and $3$, in reasonable agreement with Novak et al. (2017) and Driver et al. (2018). However, it falls short by $\approx 0.3$ dex compared to the measurements of Gruppioni et al. (2020), Enia et al. (2022), and Cochrane et al. (2023). The factor of $\approx 2$ discrepancy at $z\approx 2$ between the CSFRD computed from intrinsic SFRs in simulations and that derived from direct SFR indicators in observations, with the former being lower, has been seen in nearly all modern galaxy formation simulations in representative volumes that predict a realistic $z=0$ galaxy population, including eagle, IllustrisTNG, and Simba (Furlong et al., 2015; Pillepich et al., 2018; Davé et al., 2019). Furthermore, the same discrepancy has long been seen in observational data themselves: the SFH derived from the observationally inferred CSMD is typically a factor of $\approx 2$ lower around cosmic noon compared to the SFH constructed from observationally inferred SFRs (e.g. Wilkins et al., 2008; Yu & Wang, 2016; Wilkins et al., 2019; Shuntov et al., 2025). The ubiquity of this discrepancy was clearly demonstrated by Behroozi et al. (2019), who calibrated the semi-empirical model UniverseMachine by fitting it to the observed CSFRD and CSMD, among other constraints. Although by design the UniverseMachine is highly flexible, the authors still had to introduce a redshift-dependent correction factor to the model’s SFRs to satisfy the CSMD and CSFRD constraints simultaneously, which increased the ’observed’ SFRs relative to the true SFRs predicted by the UniverseMachine by $\approx 0.3$ dex at $z=2$. The source of this discrepancy remains unclear, but it is often attributed to systematic offsets in observational data, arising from uncertainties in the calibration of SFR indicators and assumptions regarding the dust attenuation curve, the stellar IMF, and/or gas metallicity.